OpenStudy (anonymous):

tell whether the expressions in each pair are equivalent. 6m +1 and 1 . 6 + m 9y and 9 + y mp + nq and + np -(5 - 9) and 9 - 5 8 - 4c and - 8 3(5 + z) and 15 +z 6t - 4 and 2[(2 + 1)t - 2] wwx . yz and v . w . zxy

OpenStudy (danjs):

6m + 1 = ? 1.6 + m

OpenStudy (danjs):

That first one is not equivalent, you can check by letting m be any number, for simplicity let m=0 6*0 +1 = 1.6 + 0 1 = 1.6 FALSE

OpenStudy (danjs):

What about 9*y = 9 + y those the same?